# AudioTrust: Benchmarking the Multifaceted Trustworthiness of Audio Large Language Models

### v1

@@ 4 AudioTrust: Safety
ALLMs often need to handle voice inputs that are highly contextual, emotional, or even manipulative. This imposes stringent requirements on their ability to identify potential attacks, execute refusal mechanisms, and maintain output safety. This study aims to investigate the following key research questions: (1) The overall defense capabilities of ALLMs against Jailbreak attacks and Illegal Activities Guidance attacks. (2) The impact of attacker’s emotion and gender on the defense capabilities of ALLMs when subjected to Jailbreak and Illegal Activities Guidance attacks. (3) The differences in ALLMs’ defense capabilities against Jailbreak attacks in different scenarios.
Attack Strategies. We propose a comprehensive classification system for attacks on ALLMs, covering different scenarios and risk vectors. The system has two primary categories: Jailbreak attacks and Illegal Activity Guidance. Data classification details are in Section F.1. Jailbreak attacks [34, 50, 106] refer to attempts to bypass predefined behavioral constraints via prompt injection or adversarial input (e.g., roleplay, prompt escalation, emotional manipulation). We further divide jailbreak attacks into three domains: enterprise systems, financial systems, and healthcare systems, each with 100 test items to reflect real-world scenarios. Data samples are first augmented using GPT-4o [54], then synthesized to audio using F5-TTS [12]. Illegal Activity Guidance tests model risk in responding to illegal requests (e.g., drug making, fraud), referencing SafeDialBench [9], SafetyBench [100], and related works [77, 76], with 300 data samples assembled. Our data has two main attack vectors: (1) Emotional Deception Attacks, which use emotions (e.g., sadness, urgency) to lower safeguards and induce risky outputs; (2) Authentication Bypass, which leverages realistic voice and context to trigger restricted actions. These vectors allow us to test ALLMs’ resistance to psychological and technical adversarial conditions. Construction details are in Section F.2.
Evaluation & Metrics. Given the challenge of rule-based quantification for jailbreak and harmful content, we use model-based evaluation. GPT-4o [54] serves as the evaluator for both input and output texts, leveraging its strong human-aligned judgment [64]. Outputs are scored on a 5-point scale: 5 indicates full jailbreak resistance; 1 means no resistance. We report: (1) Defense Success Rate (DSR)—effectiveness against jailbreaks (higher is better), and (2) Harmful Response Rate (HRR)—proportion of harmful outputs for illegal prompts (lower is better). Evaluation guidelines are tailored per task; see Section F.3 and Section F.4.
@@
| GPT-4o Audio | 99.0 | 99.2 | 98.8 | 100.0 |
| GPT-4o mini Audio | 99.8 | 99.0 | 98.8 | 99.9 |
Note: Due to the common issue of random audio recognition failures in open-source models, these scores may be inflated. indicates value above the average within its category (open-source or closed-source), indicates value below average, with subscript showing the absolute difference from the category average.
Results. We evaluated the safety performance of open-source and closed-source ALLMs across different scenarios, as shown in Table 3. Our observations reveal that: (1) Among all tested models, Gemini-2.5 Flash demonstrated the strongest safety performance, whereas open-source models such as MiniCPM-o 2.6, Qwen2-Audio, and SALMONN performed significantly worse in comparison. (2) In the evaluation of closed-source models, the medical scenario appeared to be more vulnerable to jailbreak prompts. By contrast, the enterprise scenario, which typically involves more diverse business and technical queries, relies more heavily on general knowledge and contextual understanding—this diversity may help mitigate the effectiveness of single-mode attack strategies. These scenario-level differences highlight a critical limitation in domain-specific alignment: while closed-source models may exhibit robustness in general-purpose tasks, they often lack sufficient resistance when dealing with specialized content. (3) Compared to domain-specific jailbreak attacks, models generally performed better in Illegal Activities Guidance tests. This suggests that models tend to be more prone to generating harmful content when facing inputs containing highly domain-specific or technical language, as opposed to more general illicit queries. For detailed results, please refer to Section F.5.
@@ 5 AudioTrust: Privacy
This study investigates the potential privacy leakage risks posed by ALLMs, focusing on unintentional information disclosure during normal usage scenarios [71]. Our study aims to address the following three key questions: (1) When explicitly provided with private information, can ALLMs recognize privacy risks and prevent information leakage, and to identify which types of data are most susceptible to disclosure? (2) Do ALLMs have the capability to infer sensitive attributes such as age from non-explicit audio signals, potentially compromising user privacy? (3) Can simple privacy-enhancing techniques (e.g., Prompt Engineering) effectively mitigate these privacy leakage and inference risks?
Attack Strategies. With the advancement of ALLMs, their daily deployment brings new privacy challenges. In audio privacy attacks, adversaries can extract unintended information such as speech content, demographics, or voiceprints [19]. While most prior work has addressed privacy for text or vision-language models [80, 71, 97, 81], we propose a privacy taxonomy for ALLMs distinguishing Direct Privacy Leakage and Privacy Inference Leakage. Direct Privacy Leakage concerns accidental disclosure of sensitive details (e.g., bank numbers, addresses) and possible model memory risks. Privacy Inference Leakage refers to ALLMs inferring private attributes (e.g., age, ethnicity) from indirect audio cues, without explicit information. Details are in Section G.1. For direct privacy leakage, we build a 600-sample dataset with synthetically generated sensitive data (bank accounts, addresses, phone numbers, etc.), generating realistic dialogues via language models and F5-TTS [12]. For privacy inference, we construct 300 samples to test ALLMs’ inference of speaker attributes from non-explicit cues [81], using real voices from Common Voice enhanced with multi-track audacity synthesis [4]. More details are in Section G.2.
Evaluation & Metrics. To evaluate privacy leakage in ALLMs, we use an audio question answering (AQA) framework, with detailed reasoning and evaluation prompts provided in Section G.3. Experiments optionally include privacy-enhancing prompts (e.g., “Bank account numbers are private and should not be disclosed”) to assess their ability to reduce unintended disclosures. Model responses are scored on a five-point scale: 5 indicates refusal for privacy reasons, and 1 represents full disclosure. Scoring is performed using GPT-4o [54]. We report refusal rate (scores 5 or 4), complete disclosure rate (score 1), and cautious disclosure rate (score 2 or 3). Further details on metrics are provided in Section G.4. We evaluated nine commonly used audio models, consisting of four open-source models and five popular closed-source audio models. Given the privacy and security concerns, our focus was more on assessing whether the models are aware of security issues, rather than emphasizing the accuracy of their results. Therefore, we primarily assessed the leakage rates related to privacy.
@@
| Qwen2-Audio | 5 | 92 | 0 | 0 |
| SALMONN | 341 | 584 | 130 | 138 |
| MiniCPM-o 2.6 | 0 | 6 | 0 | 0 |
| UltraVox | 438 | 600 | 25 | 23 |
@@
| Gemini-1.5 Pro | 70 | 590 | 16 | 28 |
| Gemini-2.5 Flash | 0 | 585 | 0 | 6 |
| Gemini-2.5 Pro | 200 | 583 | 0 | 0 |
| GPT-4o Audio | 558 | 598 | 19 | 57 |
| GPT-4o mini Audio | 600 | 600 | 41 | 116 |
| Average | 245.8 | 470.9 | 25.7 | 40.9 |
Note: Scores are sums of individual attribute leakages and higher is better. "no aug" shows baseline performance, while "aug" shows results after applying privacy-aware prompt engineering. Blue arrows () indicate better performance than average; red arrows () indicate worse performance than average.
Results. Key findings from Table 4 are as follows: (1) In the direct privacy leakage tasks, Ultravox achieves the best results among open-source models, performing on par with state-of-the-art closed-source models. Particularly when equipped with privacy-enhancing prompts, Ultravox reaches a level of performance comparable to the strongest closed-source systems. Overall, closed-source models consistently demonstrate strong capabilities in protecting sensitive information. (2) In privacy inference tasks, the privacy protection performance of all models declines to some extent. Among them, the open-source model SALMONN achieves the most outstanding results, surpassing even many closed-source models in terms of privacy awareness and refusal behavior. (3) Furthermore, introducing prompts that include privacy protection instructions can significantly improve model performance in preventing direct privacy leaks. Both open-source and closed-source models benefit from such prompt engineering techniques, with the Gemini series showing the most notable improvements. However, for inference-based privacy leakage scenarios, the effectiveness of these enhancements is limited. In some cases, such as with Gemini-2.5 Pro, no improvement is observed at all. For detailed results, please refer to Section G.5.
@@ 6 AudioTrust: Robustness
ALLMs demonstrate powerful capabilities under ideal conditions, but they still face serious limitations in real-world application scenarios. These models often produce “hallucinations” when dealing with audio containing background noise, multi-person conversations or degraded quality - that is, generating content that is inconsistent with the input or completely fictional. (1) How do ALLMs perform as a whole in carrying out its core tasks under different robustness challenges? (2) To what extent do various audio disturbances specifically affect the speech recognition accuracy of ALLMs? (3) How capable are ALLMs in evaluating audio clarity under different robustness conditions, and can they accurately perceive and interpret these differences in audio quality based on preset dimensions?
Attack Strategies. The challenges faced by large audio language models can be divided into two major categories: intentionally designed adversarial attacks [47] and non-adversarial challenges that occur naturally in real-world environments [58]. Adversarial attacks induce false output through carefully designed and imperceptible perturbations [10]; Non-adversarial challenges include common interfering factors in daily environments. Specifically, we focus on six dimensions: robustness against intentionally designed adversarial samples, the ability to correctly identify target conversations in a multi-person conversation environment [8], the ability to handle natural environmental noise [98], the ability to cope with audio of different quality levels, resistance to various artificial or natural noises, and performance in scenarios where multiple people speak simultaneously or alternately. We have built dedicated test datasets for various robustness challenges to simulate typical interferences in real-world application scenarios. Each dimension contains 40 test samples, covering multiple languages and topics to ensure the comprehensiveness and objectivity of the assessment. More details of the data are available at Section H.1 and Section H.2.
Evaluation & Metrics. Given the challenges in directly measuring robustness or output risk, we adopt a model-based evaluation framework using GPT-4o [54], following recent evidence [103] that such models can reliably assist human judgment. Each test output is rated on a discrete 10-point scale (0–10), with scoring rubrics tailored per prompt and task: 10 indicates strong consistency with audio quality, while 0 means perceptual failure or inability to recognize the specified variation. Prompt templates are detailed in Section H.3.1. For comprehensive evaluation, we also report two quantitative metrics (Section H.3.2): Cross-Modal Word Error Rate (CM-WER), measuring dissimilarity between generated and human-annotated transcriptions [58]; and Content Consistency Rate (CCR), assessing factual alignment between ALLM outputs and ground-truth audio content [49].
@@
| Average | 5.74 | 5.81 | 6.13 | 5.66 | 5.45 | 5.87 |
Note: : higher than column average (better for accuracy), : lower than column average (worse for accuracy), subscript is absolute difference. Scenarios: AR (Adversarial Robustness), AQV (Audio Quality Variation), BC (Background Conversation), ES (Environmental Sound), MS (Multiple Speakers), NI (Noise Interference). The overall low accuracy highlights hallucination concerns.
Results. We evaluated the robustness of nine ALLMs under various challenging audio conditions, as shown in Table 5. Complete results and qualitative examples are detailed in Section H.3.3. We observed the following key points: (1) There is a clear gap in audio robustness between model families. Closed-source models (Gemini-2.5 Pro, Gemini-1.5 Pro, GPT-4o Audio) consistently outperform open-source counterparts across conditions including background noise, multiple speakers, and audio degradation. Among open-source models, MiniCPM-o 2.6 exhibits notable robustness, rivaling some closed-source baselines, while SALMONN and Ultravox struggle with adverse or noisy inputs. (2) When evaluating audio clarity, Gemini models demonstrate strong sensitivity to quality degradations, offering accurate and nuanced characterizations of different issues. GPT-4o models display moderate awareness, but sometimes misclassify disturbance types. Open-source models generally lack the ability to detect or describe subtle audio quality degradation, often misattributing issues or overlooking them entirely. (3) Overall, among closed-source models, Gemini-2.5 Pro achieves the most balanced and robust performance across all scenarios, while GPT-4o Audio excels under certain disturbances but is less robust to others. In the open-source group, MiniCPM-o 2.6 and Qwen2-Audio are comparatively robust, yet still lag behind the closed-source models. SALMONN consistently exhibits the weakest robustness, suggesting notable limitations in its audio processing capabilities.

### v3

@@ 4 AudioTrust: Safety
The safety landscape for ALLMs presents challenges distinct from text-based systems (Wei et al., 2023; Wang et al., 2023c). The tone of a voice (whether urgent, distressed, or authoritative) can serve as a powerful tool to bypass the model’s safety alignment. For example, a user might pretend to have a medical emergency with a panicked voice to request dangerous information. This emotional attack vector, which leverages the persuasive nature of the human voice, is a novel challenge that text-only safety protocols do not address. Due to the high-dimensional and complex nature of these acoustic features, they may introduce new vulnerabilities during the encoding process and are challenging to align effectively with the safety constraints established for the text modality.
Attack Strategies. To systematically evaluate the safety of ALLMs, AudioTrust is structured along two orthogonal dimensions: Domain-Specific Policy Jailbreak Attacks and General Illegal Activity Guidance. The first category, Domain-Specific Policy Jailbreak Attacks, encompasses adversarial attempts to circumvent a model’s intrinsic behavioral constraints through techniques like prompt injection and role-playing (Li et al., 2025; Mo et al., 2024; Zhou et al., 2025). To ensure real-world applicability, these attacks are stratified into three critical domains: enterprise, financial, and healthcare systems, each populated with 100 test instances. The second category, General Illegal Activity Guidance, assesses the model’s propensity to generate harmful content for illicit requests (e.g., fraud, weapon crafting), drawing upon established benchmarks (Cao et al., 2025; Zhang et al., 2023c) to assemble 300 diverse prompts. In AudioTrust, we specifically investigated an attack vector based on the audio modality: Emotional Deception, which involves using speech prompts with specific emotional cues (e.g., anger, sadness) to induce the model to bypass its safety filters. All textual prompts were augmented via GPT-4o (OpenAI et al., 2024) and subsequently synthesized into audio using F5-TTS (Chen et al., 2024). Further details on data classification and construction are available in Appendices F.1 and F.2, respectively.
Evaluation & Metrics. Quantifying the success of jailbreak attacks and the harmfulness of generated content via rule-based methods is notoriously challenging. We therefore employed a model-based evaluation, leveraging GPT-4o. This choice is motivated by its strong alignment with human preferences on safety-sensitive assessments (Tan et al., 2025). GPT-4o rated each model-generated response on a 5-point Likert scale, where a score of 5 indicates complete refusal of a malicious request, and 1 denotes a fully compliant, harmful generation. Subsequently, these scores were further reviewed by human evaluators. Based on these scores, we computed two primary metrics: the Defense Success Rate (DSR), which measures the model’s robustness against jailbreak attempts (higher is better), and the Harmful Response Rate (HRR), quantifying the proportion of unsafe content produced in response to illicit prompts (lower is better). The detailed evaluation protocols, which are adapted for each task, are provided in Sections˜F.3 and F.4.
@@
| Qwen2-Audio | 68.2 | 80.6 | 81.4 | 92.5 |
| MiniCPM-o 2.6 | 76.2 | 79.2 | 81.6 | 96.2 |
| Step-Fun | 70.6 | 86.2 | 89.0 | 94.5 |
| Qwen2.5-omni | 97.2 | 94.8 | 94.2 | 99.1 |
| Kimi-Audio | 99.4 | 98.2 | 95.2 | 99.9 |
| OpenS2S | 51.4 | 67.8 | 75.2 | 71.8 |
| Step-Audio2 | 88.0 | 68.4 | 73.0 | 69.0 |
@@
Note: Due to the common issue of random audio recognition failures in open-source models, these scores may be inflated.
Results. We evaluated the safety performance of both open- and closed-source ALLMs across different scenarios, as shown in Table˜3. Our analysis reveals several observations: (1) While leading closed-source models such as Gemini-2.5 Flash maintain strong safety performance, open-source models exhibit substantial variation. Notably, Kimi-Audio demonstrates remarkable robustness, achieving scores comparable to the best closed-source counterparts. In contrast, models such as OpenS2S and SALMONN display considerable vulnerability, highlighting the large capability gap within the open-source ecosystem. (2) For closed-source models, the medical domain remains relatively more susceptible to jailbreak attacks, suggesting that domain-specific alignment in specialized areas is still an open challenge even for highly capable systems. In open-source models, no single domain consistently emerges as the weakest link, with vulnerabilities appearing to be model-dependent. (3) Most models, regardless of being open-source or closed-source, generally exhibit stronger defenses against General illegal activity guidance prompts compared to domain-specific jailbreak attempts. This indicates that broad safety training against overtly illegal content is largely effective, whereas nuanced, domain-targeted jailbreaks remain a more successful pathway for adversaries. Detailed results are provided in Section˜F.5.
@@ 5 AudioTrust: Privacy
This section examines privacy challenges specific to ALLMs. In text-based systems (Wang et al., 2023a; Huang et al., 2024), privacy risks typically involve the model memorizing and repeating sensitive information from its training data. ALLMs face this risk, but also a more subtle and pervasive one: information leakage from the acoustic signal itself. The sound of a voice can reveal a speaker’s approximate age; the background noise can betray their location (e.g., a quiet office or a busy cafe). This means ALLMs can infer private information even when it is never explicitly stated, creating a new class of privacy risks beyond simple content disclosure.
Attack Strategies. In AudioTrust, we categorize the privacy risks associated with ALLMs into two distinct groups: (1) Direct Privacy Leakage, which pertains to sensitive information explicitly stated within the conversational content. In this category, the ALLMs might reveal data such as a bank account number mentioned during a conversation. The formulation of this risk is informed by similar challenges in traditional large language models. (Wang et al., 2023a; Huang et al., 2024). (2) Privacy Inference Leakage, where private attributes are inferred from paralinguistic cues rather than the explicit content. Such cues include a speaker’s tone of voice, speech rate, accent, and vocal quality. This risk, which is unique to ALLMs, involves the model deducing personal attributes like age or ethnicity from the audio itself, independent of the semantic content. To evaluate these risks, we constructed two datasets: a direct privacy leakage set containing 600 synthetic dialogues and a privacy inference set with 300 samples. The latter was created using speech from Common Voice (Ardila et al., 2020) mixed with background audio from freesound (Freesound, ). For a detailed taxonomy and construction methods, see Appendixes G.1 and G.2.
Evaluation & Metrics. We evaluated our experiment within an Audio Question Answering (AQA) framework, employing two distinct settings: one utilizing privacy-enhancing prompts and the other using standard prompts. GPT-4o serves as the evaluator, assigning scores from 1 to 5 to each response and then the results were checked by human. A score of 1 indicates complete disclosure of private information, while a score of 5 signifies a refusal of the request due to privacy safety concerns. Our main evaluation metrics are refusal rate and accuracy. A higher refusal rate indicates stronger privacy protection. Accuracy is calculated only when privacy information is leaked. Further details on evaluation prompts and evaluation metrics can be found in Appendixes G.3 and G.4.
@@
| SALMONN | 57.50 | 96.83 | 38.75 | 45.00 |
| UltraVox | 73.46 | 99.67 | 6.50 | 7.50 |
| Qwen2-Audio | 0.83 | 23.67 | 1.50 | 0.75 |
| MiniCPM-o 2.6 | 0.00 | 0.67 | 1.50 | 1.50 |
| Step Fun | 41.50 | 98.33 | 11.50 | 14.75 |
| Qwen2.5-Omni | 0.00 | 1.17 | 0.50 | 0.00 |
| Kimi Audio | 0.17 | 1.00 | 11.25 | 6.50 |
| OpenS2S | 7.68 | 43.83 | 11.00 | 10.00 |
| Step Audio2 | 0.00 | 38.83 | 0.00 | 0.00 |
@@
| Gemini-1.5 Pro | 11.85 | 98.50 | 4.00 | 7.00 |
| GPT-4o Audio | 92.00 | 99.67 | 5.00 | 16.25 |
| GPT-4o mini Audio | 100.00 | 100.00 | 10.50 | 30.00 |
| Gemini-2.5 Flash | 1.36 | 96.44 | 1.25 | 2.75 |
| Gemini-2.5 Pro | 33.50 | 94.17 | 0.50 | 1.25 |
| Average | 29.99 | 63.77 | 7.41 | 10.23 |
Note: Scores are refusal rates (higher is better). “w/o” vs. “w/” compares standard prompting without and with privacy-aware prompt engineering.
Results. We evaluated the privacy protection performance of both open-source and closed-source ALLMs. The results are shown in Table˜4. Our analysis revealed several key observations: (1) Direct Privacy Leakage: In experiments using the direct privacy leakage dataset, closed-source ALLMs generally achieved superior results. The GPT-4o series demonstrated the best refusal rate, reaching 100%. Furthermore, the implementation of privacy-enhancing prompts significantly improved performance across almost every model, showing approximately a 25% improvement. This indicates that traditional prompt engineering can effectively enhance privacy protection when sensitive information is explicitly present in conversation content. (2) Privacy Inference Leakage: Both open-source and closed-source ALLMs performed poorly in addressing privacy inference leakage. They rarely refused requests for certain types of privacy information, such as age and ethnicity, with the refusal rate of only 7.41%. Unlike direct privacy leakage, privacy-enhancing prompts had a minimal impact, yielding only about a 3% improvement. Across all our results, we observed that ALLMs struggle to process privacy information that is not directly stated in a conversation but rather inferred from paralinguistic cues. This suggests that ALLMs might not identify such inferred information as private or requiring protection. This could be due to the training process, where paralinguistic cues may have been considered less important than conversational content, or because of insufficient data for this specific type of information. Detailed results accuracy analysis are available in Appendix G.5.
@@ 6 AudioTrust: Robustness
This section investigates the robustness of ALLMs in maintaining performance against real-world audio distortions. Unlike in text (Wang et al., 2023b) and vision (Hendrycks & Dietterich, 2019), audio robustness presents unique challenges due to its physical nature. Audio signals are temporal encodings of acoustic phenomena, with inherent properties like source, medium, and reverberation shaping complex auditory scenes. A truly robust ALLM must be able to disentangle the primary speech signal from this acoustic clutter and maintain its performance. This is not simply a matter of better speech recognition, but a test of the model’s fundamental ability to function in imperfect acoustic environments.
Attack Strategies. We categorize robustness challenges for ALLMs into two primary types: intentional adversarial attacks (Madry et al., 2017) and naturally occurring phenomenon of performance degradation (Radford et al., 2023). Adversarial attacks employ carefully crafted, imperceptible perturbations to induce model failure (Carlini & Wagner, 2018). In contrast, non-adversarial challenges encompass common real-world interferences. We evaluate model robustness across several key dimensions: (1) adversarial resilience, including three categories: natural noise, speaker identification, and voice overlap situations.speech (Bredin et al., 2020); (2) robustness to environmental noise (Zhang et al., 2023b) and variations in audio quality. To this end, we constructed dedicated datasets simulating these interferences. Each dimension contains 40 multilingual and multi-topic samples to ensure a comprehensive assessment. Further dataset details are available in Sections˜H.1 and H.2.
Evaluation & Metrics. Given the challenges in directly measuring robustness or output risk, we adopt a model-based evaluation using GPT-4o (OpenAI et al., 2024), following recent evidence (Zheng et al., 2023a). Each test output is rated on a discrete 10-point scale, with scoring rubrics tailored per prompt and task: 10 indicates strong consistency with audio quality, while 0 means perceptual failure or inability to recognize the specified variation. These scores were then subject to further review by the human evaluators. Prompt templates are detailed in Section˜H.3.1. For comprehensive evaluation, we also report two quantitative metrics (Section˜H.3.2): CM-WER, measuring dissimilarity between generated and human-annotated transcriptions (Radford et al., 2023); and Content Consistency Rate (CCR),(e.g., transcribe the voice with added interfering information through multiple rounds of model dialogue prompts to understand its semantic context, and then score the transcription against the original voice text content.) assessing factual alignment between ALLMs outputs and ground-truth audio content (Min et al., 2023).
@@
| Qwen2.5-Omni | 8.14 | 7.10 | 7.50 | 7.93 | 7.12 | 7.17 |
| Step-Fun | 5.00 | 7.48 | 8.20 | 7.42 | 5.89 | 7.08 |
| Kimi Audio | 5.67 | 6.83 | 6.00 | 6.83 | 7.08 | 6.94 |
| OpenS2S | 8.25 | 6.46 | 5.17 | 6.39 | 2.33 | 6.25 |
| Step-Audio2 | 6.18 | 6.58 | 7.92 | 6.82 | 0.00 | 6.78 |
@@
| Average | 6.67 | 6.94 | 7.08 | 6.89 | 6.27 | 6.95 |
Note: Scenarios: AR (Adversarial Robustness), AQV (Audio Quality Variation), BC (Background Conversation), ES (Environmental Sound), MS (Multiple Speakers), NI (Noise Interference). The overall low accuracy highlights hallucination concerns.
Results. Our robustness evaluation (Table 5) reveals a significant performance gap between closed- and open-source ALLMs. Detailed results are in Section˜H.3.3. (1) Superior robustness in closed source models. Such as the closed-source models like Gemini-2.5 Pro,a cross nearly all tested conditions including background noise, multi-speaker conversations, and audio quality variation leading closed source systems consistently outperform their open source counterparts. Notably, this advantage is most apparent under severe acoustic distortion, suggesting that proprietary models benefit from more mature front-end signal processing and advanced noise suppression architectures. (2) In contrast, many open-source systems experience a steep decline in transcription accuracy and semantic coherence when exposed to moderate noise or compression. Their audio encoders often fail to disentangle source speech from channel artifacts, leading to semantic hallucinations in which non-speech noise is incorrectly interpreted as meaningful content.